OpenStudy (jkbo):

Please help me Given that a linear equation has a slope of 2, and a y-intercept of 8, determine the slope-intercept and standard forms of the line. A.y = 2x + 8; 2x + y = 8 B.y = –2x + 8; 2x – y = –8 C.y = 2x + 8; 2x – y = –8 D.y = –2x + 8; 2x + y = 8

jimthompson5910 (jim_thompson5910):

hint: slope-intercept form is y = mx+b m = slope b = y-intercept

OpenStudy (jkbo):

ok so It would be A. then @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are y = 2x + 8 and 2x + y = 8 equivalent?

OpenStudy (jkbo):

No but its the only thing positive @jim_thompson5910

jimthompson5910 (jim_thompson5910):

solve 2x + y = 8 for y to get _________

OpenStudy (jkbo):

I got y=8-2x @jim_thompson5910

jimthompson5910 (jim_thompson5910):

which is the same as y = -2x + 8

jimthompson5910 (jim_thompson5910):

so clearly y = 2x + 8 and 2x + y = 8 are NOT equivalent

jimthompson5910 (jim_thompson5910):

that's why choice A is false

OpenStudy (jkbo):

ok so it would be D than @jim_thompson5910

OpenStudy (jkbo):

@jim_thompson5910 ??

jimthompson5910 (jim_thompson5910):

no it's C.y = 2x + 8; 2x – y = –8

OpenStudy (jkbo):

o ok thank you